OpenStudy (blackbird02):

Divide 40 into two parts such that the sum of the squares of these parts is a minimum.

OpenStudy (gorv):

hmmm..let one part is x

OpenStudy (blackbird02):

ok, what to do next?

OpenStudy (anonymous):

\[x^2+y^2\] x = equals one side however the otherside = 40 - x \[x^2+(x-4)^2 = y\]

OpenStudy (anonymous):

wait lol \[x^2+(40-x)^2= y\]

OpenStudy (gorv):

now use differentiation to find least or max value

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

I have the answer but i need to see you're working towards getting the answer yourself

OpenStudy (blackbird02):

so it's gonna be 2x^2+1600-80x=y

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

how'd you get 1600 - 80 x

OpenStudy (anonymous):

derivative rule \[x^n = nx \]

OpenStudy (blackbird02):

i simplified (40-x)^2

OpenStudy (anonymous):

and the other is \[(x+c)^n=n(x+c)(x+c)'\]

OpenStudy (anonymous):

ehh you can do it that way but if you just use the rule that i showed you... it comes out to be really simple

OpenStudy (anonymous):

\[2x +2(40-x)(-1) = y'\] \[2x+(80-2x)(-1) = 2x - 80 +2x = 4x-80\] \[4x - 80 = y '\]

OpenStudy (anonymous):

but it looks right if you factor out everything

OpenStudy (anonymous):

so you just take the derivative of that and you'll get what i got, correct?

OpenStudy (anonymous):

simple rules ... you only need the first rule when factoring out.... anyways, lastly, what does your book say about y' = 0

OpenStudy (anonymous):

@blackbird02

OpenStudy (blackbird02):

there wasn't anything stated about it

OpenStudy (anonymous):

alright well think of it this way... y' or your derivative is the slope of your equation correct?

OpenStudy (anonymous):

well y' is a tangent line with the same slope as your line... therefore if you have say...... x^2 as your equation...

OpenStudy (anonymous):